# A tuneable frequency comb via dual-beam laser-solid harmonic generation

### v1

@@ I Introduction
In “classic” high harmonic generation (HHG) by high-intensity laser beams in isotropic configurations, the frequency spectrum contains odd integer multiples of the frequency of the pump beam, e.g. for HHG in ionised gas ferray or HHG via normal incidence onto a solid target lichters . The wavelength of hte pump laser ranges from the mid- and near-infrared (IR) to the ultraviolet, which in turn changes the HHG spectrum from a coherent supercontinuum to a collection of narrowband peaks popmin12 ; popmin15 . However, greater control of the frequency spectrum, especially larger spacing between the peaks, is important for numerous applications, e.g. to tune the effective carrier frequency of the attosecond pulses created via HHG. Hence the concept of the “frequency comb”, where only a subset of peaks from the harmonic spectrum is selected, not all peaks, see e.g. alon98 ; bayku21 ; rego22 ; trines24 . The specific subset of selected peaks is tuneable via higher order properties (transverse shape, polarisation, etc.) of either the driving beams or the target used in the HHG process.
@@
Recently, Rego et al. rego22 developed the concept of the “necklace beam”. In this configuration, two beams with parallel linear polarisations and equal frequencies but different orbital angular momentum (OAM) levels, no common divisor and not both odd, e.g. $\ell_{A}=+2$ and $\ell_{B}=-3$, are used to drive harmonics in an isotropic medium (gas). When the harmonic light is observed through an on-axis pinhole, only the light with zero OAM is caught; this light is found to contain only the harmonics $(2n+1)|\ell_{A}-\ell_{B}|$, so odd multiples of the frequency $|\ell_{A}-\ell_{B}|\omega_{0}$ rather than $\omega_{0}$. This is of course tuneable via $\ell_{A,B}$.
@@
Asymmetric frequency combs have been considered using a single CP pump laser interacting with a target with a finite rotational symmetry, see e.g. Refs. alon98 ; bayku21 ; trines24 . Again, the on-axis harmonic light is collected through a pinhole and only those harmonic multiples that have an OAM-free component are caught.
@@
In this paper, we show that the role of the OAM level $\ell$ in the “necklace beam” configuration can also be assumed by a small transverse wave vector component $k_{\perp}$. To this end, we consider two identical LP beams with frequency $\omega_{0}$ and p-polarisation, crossing at a small angle. This configuration is similar to that displayed in Figure 3a of Hickstein et al. hickstein . However, we turn the question around: rather than asking “for a given frequency, which emission angles do we see”, we ask “for a given emission angle, which frequencies do we see”. We predict that for a beam crossing half-angle $\alpha_{0}$ and a viewing angle $\alpha_{v}$ with $\tan\alpha_{0}=N\tan\alpha_{v}\ (N\in\mathbb{N})$, the harmonic spectrum contains peaks at $(2n+1)N\omega_{0}$. Thus, $N$ takes the role of $|\ell_{A}-\ell_{B}|$ in the necklace beam configuration. We note that $N$ is tuned via the observation angle $\alpha_{1}$, which is easier and more versatile than tuning via $|\ell_{A}-\ell_{B}|$. It also allows for more extensions, which will be discussed below.
@@ II Theory
For a scenario with two identical beams crossing at a shallow angle hickstein , the harmonic progression reads:
@@ III Simulation setup
We carried out 2-D particle-in-cell simulations using the Epoch code epoch1 ; epoch2 to model the interaction of two laser beams hitting a flat target at near-normal incidence. Numerical parameters were: a box size of $16\times 64\ \mu\mathrm{m}^{2}$, a grid size of $4096\times 8192$ cells, 10 particles per cell, and a total simulation time of 250 fs. Physical parameters were: a laser wave length of 1 $\mu$m, a laser peak intensity of $3\cdot 10^{19}$ W/cm2, a beam waist of 16 $\mu$m, a transverse pulse envelope that is either Gaussian ($\exp(-r^{2})$) or super-Gaussian ($\exp(-r^{16})$), a Gaussian longitudinal envelope with a duration of 50 fs, a plasma slab with a maximum density of $8n_{\mathrm{cr}}$ and a pre-plasma layer extending 1 $\mu$m in front of the slab. The laser beams are identical and impact the same spot at angles of ±$\pm 15$ degrees with respect to the target normal.

### v2

@@ I Introduction
In “classic” high harmonic generation (HHG) by high-intensity laser beams in isotropic configurations, the frequency spectrum contains odd integer multiples of the frequency of the pump beam, e.g. for HHG in ionised gas ferray or HHG via normal incidence onto a solid target lichters . The wavelength of hte pump laser ranges from the mid- and near-infrared (IR) to the ultraviolet, which in turn changes the HHG spectrum from a coherent supercontinuum to a collection of narrowband peaks popmin12 ; popmin15 . However, greater control of the frequency spectrum, especially larger spacing between the peaks, is important for numerous applications, e.g. to tune the effective carrier frequency of the attosecond pulses created via HHG. Hence the concept of the “frequency comb”, where only a subset of peaks from the harmonic spectrum is selected, not all peaks, see e.g. alon98 ; bayku21 ; rego22 ; trines24 . The specific subset of selected peaks is tuneable via higher order properties (e.g. transverse shape, polarisation or phase) of either the driving beams or the target used in the HHG process.
@@
Recently, Rego et al. rego22 developed the concept of the “necklace beam”. In this configuration, two beams with parallel linear polarisations (LP) and equal frequencies but different orbital angular momentum (OAM) levels, no common divisor and not both odd, e.g. $\ell_{A}=+2$ and $\ell_{B}=-3$, are used to drive harmonics in an isotropic medium (gas). When the harmonic light is observed through an on-axis pinhole, only the light with zero OAM is caught; this light is found to contain only the harmonics $(2n+1)|\ell_{A}-\ell_{B}|$, so odd multiples of the frequency $|\ell_{A}-\ell_{B}|\omega_{0}$ rather than $\omega_{0}$. This is of course tuneable via $\ell_{A,B}$.
@@
Asymmetric frequency combs have been considered using a single pump laser with circular polarisation (CP) interacting with a target with a finite rotational symmetry, see e.g. Refs. alon98 ; bayku21 ; trines24 . Again, the on-axis harmonic light is collected through a pinhole and only those harmonic multiples that have an OAM-free component are caught.
@@
In this paper, we show that the role of the OAM level $\ell$ in the “necklace beam” configuration can also be assumed by a small transverse wave vector component $k_{\perp}$. To this end, we consider two identical LP beams with frequency $\omega_{0}$ and s-polarisation, crossing at a small angle. This configuration is similar to that displayed in Figure 3a of Hickstein et al. hickstein . However, we turn the question around: rather than asking “for a given frequency, which emission angles do we see”, we ask “for a given emission angle, which frequencies do we see”. We predict that for a beam crossing half-angle $\alpha_{0}$ and a viewing angle $\alpha_{v}$ with $\tan\alpha_{0}=N\tan\alpha_{v}\ (N\in\mathbb{N})$, the harmonic spectrum contains peaks at $(2n+1)N\omega_{0}$. Thus, $N$ takes the role of $|\ell_{A}-\ell_{B}|$ in the necklace beam configuration. We note that $N$ is tuned via the observation angle $\alpha_{1}$, which is easier and more versatile than tuning via $|\ell_{A}-\ell_{B}|$. It also allows for more extensions, which will be discussed below.
@@ II Theory
For a scenario with two identical s-polarised LP beams crossing at a shallow angle, as used by e.g. Hickstein et al. hickstein , the harmonic progression reads:
@@ III Simulation setup
We carried out 2-D particle-in-cell simulations using the Epoch code epoch1 ; epoch2 to model the interaction of two laser beams hitting a flat target at near-normal incidence. Numerical parameters were: a box size of $16\times 64\ \mu\mathrm{m}^{2}$, a grid size of $4096\times 8192$ cells, 10 particles per cell, and a total simulation time of 250 fs. Physical parameters were: a laser wave length of 1 $\mu$m, a laser peak intensity of $3\cdot 10^{19}$ W/cm2, a beam waist of 16 $\mu$m, a transverse pulse envelope that is either Gaussian ($\exp(-r^{2})$) or super-Gaussian ($\exp(-r^{16})$), a Gaussian longitudinal envelope with a duration of 50 fs, a plasma slab with a maximum density of $8n_{\mathrm{cr}}$ and a pre-plasma layer extending 1 $\mu$m in front of the slab. The laser beams are identical with linear s-polarisation and impact the same spot at angles of ±$\pm 15$ degrees with respect to the target normal.
@@ V Extensions
  6. 6.
    When the harmonic progression has more than 2 dimensions, e.g. 3-D, then one can either choose a 2-D subset on a plane or a 1-D subset on a line. This offers an extra degree of freedom in selecting frequency combs and thus deserves further exploration.